OpenStudy (anonymous):

Will medal and fan Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received: x − 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values. Numerical Answers Expected!

OpenStudy (texaschic101):

y = total number of votes cast = 125000 so sub that in.. x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 --- add 62,500 to both sides x = 31,250 + 62,500 x = 93750